# The differences of baryonic and dark matter scaling relations in galaxy clusters between IllustrisTNG simulations and observations

### v1

@@ 4.2 Looking for the missing baryons
A limitation usually found in observations is the lack of information on the cold and warm gas components. As simulations do not have these issues, we will estimate which fraction of baryons we are undetecting when only considering the hot gas from the ICM. We divided the baryonic matter into cold gas, warm gas, and hot gas for the gaseous component; and into galaxy stellar mass (from $r_{\rm op}$ and subfind) and intra-cluster light (ICL) for the stellar component. Additionally, we split our clusters between low-mass $\log_{10}{(M_{\rm 500c}/{\rm M_{\odot}})}<14$ and high-mass haloes $\log_{10}{(M_{\rm 500c}/{\rm M_{\odot}})}\geq 14$. Our sample comprises 137 and 1115 low-mass haloes in IllustrisTNG-100 and IllustrisTNG-300, and 81 and 490 high-mass haloes in IllustrisTNG-100 and IllustrisTNG-300, respectively. Additionally, for visualization purposes, we include the complete sample (labeled as ”all halo types”) in the following figures.
@@
Additionally, Table 3 shows the average baryonic mass values for components not considered in C18 observations: cold gas, warm gas, and ICL ($r_{\rm op}$). The sum of those components is defined as the missing baryons from these simulations. The data in Table 3 reveal a tendency for low-mass haloes to contain more missing (undetected) baryons than high-mass haloes. Specifically, low-mass haloes exhibit approximately $17.83$ per cent in TNG100 and $17.67$ per cent in TNG300 of missing baryons. In contrast, high-mass haloes have about $6.95$ per cent in TNG100 and $5.90$ per cent in TNG300 of missing baryons. In both cases, hot gas remains the most significant baryonic component within the haloes. However, warm gas and ICL ($r_{\rm op}$) correspond to a non-negligible fraction of baryons ”missed” in low-mass haloes. Conversely, most of the missing baryons are attributed to the ICL ($r_{\rm op}$) in high-mass haloes, which shows a similar proportional contribution to that observed in low-mass haloes. Finally, Table 3 also presents the stellar components measured using the subfind aperture, highlighting the differences between this method and the $r_{\rm op}$ aperture in terms of stellar mass measurements. The results show that the subfind algorithm collects more stellar mass within galaxies compared to the $r_{\rm op}$ spherical aperture. This is a consequence of the subfind algorithm, which, by construction, assigns to the central galaxy all particles within the virial radius that are not associated with any satellite subhalo (Springel et al., 2001; Dolag et al., 2009). As a result, diffuse stellar components, such as the stellar halo, are included as part of the central galaxy mass.
When considering all haloes, Table 3 shows that TNG100 and TNG300 have 13.79 per cent and 14.08 per cent of missing baryons respectively. The hot gas component remains the most significant baryonic component. The overall fractions are similar to those of low-mass haloes but with lower contributions from the warm gas and ICL ($r_{\rm op}$) components, as high-mass haloes are included in this analysis. Among all components, warm gas is the most affected when considering all haloes, showing a reduction of approximately 3 per cent compared to the low-mass haloes in both simulations. The ICL ($r_{\rm op}$) component also exhibits a reduction of about $0.40$ per cent relative to the low-mass haloes in both simulations. This analysis highlights the importance of dividing the halo sample since baryons could be present in different states in high- and low-mass clusters.
@@
Gouin et al. (2022) studied the distribution of gas components in TNG300 haloes at $z=0$. These clusters were categorised by mass into clusters ($M_{\rm 200c}/{\rm{M_{\odot}}}/h>1\times 10^{14}$) and groups ($5\times 10^{13}<M_{\rm 200c}/{\rm{M_{\odot}}}/h<1\times 10^{14}$) of galaxies. Additionally, they divided the gas into five independent phases: hot gas, warm dense gas (WCGM), warm diffuse gas (WHIM), cold dense gas (halo gas), and cold diffuse gas (IGM). They calculated the gas fraction of these phases as a function of $\rm R_{200}$. Table 4 shows the relative gas percentages for our haloes, calculated based on the percentages in Table 3. Despite differences between our sample and the sample from Gouin et al. (2022), the gas fractions are remarkably similar at $R\sim 0.6R_{200}\sim R_{500}$. In galaxy groups, most of the contribution to these percentages, besides the hot gas, comes from warm diffuse gas, which corresponds to the circumgalactic medium (CGM) (Gouin et al., 2022). At this radius, WCGM gas accounts for a larger fraction than WHIM gas. In galaxy clusters, the majority of the gas fraction comes from hot gas. Based on this analysis, the gas phase associated with the missing baryons in our sample is likely dominated by WCGM gas.
These results suggest that the break in the power-law seen in Fig. 3 are driven by the higher fractions of ICL, cold gas, and warm gas present in low-mass clusters compared to massive ones. Additionally, cold gas percentages are lower in all cases compared to warm gas. Therefore, the warm gas and the ICL in low-mass haloes are the primary contributors to the observed changes in baryonic scaling relations. This interpretation aligns with the WHIM proposal, which suggests that the missing baryons identified in C18 could reside in the warm diffuse phase. Furthermore, A23 propose that missing baryons are redistributed by non-gravitational processes (e.g., AGN, stellar, and UV feedback) into a region between $1.5-2.5R_{\rm 200c}$, where the baryon fraction reaches the universal value. Both scenarios are compatible, suggesting a coherent explanation for the missing baryons problem. This conclusion could have significant implications for future observational research related to the baryon distribution, mass measurements, and calibration of scaling relations.
@@ 4.3 FoF masses in baryonic scaling relations
The changes in baryonic scaling relations when considering the total mass of haloes within a sphere radius $R_{\rm 500c}$ are explored in Subsections 4.1 and 4.2. The $R_{\rm 500c}$ is particularly useful for measuring gas masses, using X-ray and hot-the SZ effect in an observational context. In this section, we shift the mass variables from $R_{\rm 500c}$ to masses defined by the FoF algorithm. i.e., accounting for all the mass ”bound” to the cluster without a fixed aperture. This approach compares how scaling relations change when using these broader mass definitions.
@@
Figure 4: The Simple Power Law (SPL) fit (green region) of the IllustrisTNG haloes (black dots), to the scaling relation of the mass fraction of baryons $f_{\rm bar}$ and halo mass $M$ measured by friend-of-friends (FoF) algorithm, for all redshifts. All panels have their respective simulation tag and a value for the $\alpha_{\rm SPL}$ slope from the SPL fit in the upper-left corner. The fraction of baryons includes all the gas and all the stars inside the halo. The yellow region represent the Planck Collaboration et al. (2016a) fraction of baryons. All the regions are in $3\sigma$.
@@
Fig. 4 presents the SPL scaling relation models for baryonic mass fractions of IllustrisTNG haloes in the FoF delimited mass context. Comparing these results with the all-gas + all-stars scaling relations at $R_{\rm 500c}$ shown in Fig. 3, we observe a mean slope difference of $\alpha_{\rm SPL,FoF}-\alpha_{\rm SPL,500c}\simeq-0.06$ in TNG100 and $\alpha_{\rm SPL,FoF}-\alpha_{\rm SPL,500c}\simeq 0.08$ in TNG300. In the FoF context, low-mass haloes boost their baryon fractions, displaying a tendency towards $\alpha_{\rm SPL,FoF}\to 0$ slopes in both simulations. This behavior accounts for the differences observed with the $R_{\rm 500c}$ mass context. Additionally, the self-similarity of the baryon fraction is satisfied, as the $\alpha_{\rm SPL,FoF}\to 0$ results are consistent with Eq. (6). Furthermore, these slope tendencies indicate that high-mass haloes can form through the assembly of low-mass haloes, reflecting the hierarchical formation scenario. However, most IllustrisTNG haloes fail to reach the baryon fraction measured by Planck Collaboration et al. (2016a) within the $3\sigma_{\rm Planck}$ region (with $\sigma_{\rm Planck}=0.004$ being the error of the baryon fraction in Planck Collaboration et al., 2016a). Most haloes exhibit baryon fractions approximately $0.1$ dex below the mean value, with only a few exceptions. This result suggests the re-emergence of the missing baryon problem in the FoF context, which will be further discussed in Section 5.
4.4 Fitting power law models as scaling relations
@@
In Fig. 3, we observe how the SPL model fits the scaling relations of baryon fractions in IllustrisTNG haloes. This provides a framework to compare the haloes with the scaling relations from C18. However, the distribution of scattered data and the SPL model’s fit do not appear to align well. In the hot-gas + $r_{\rm op}$ star comparison, the data exhibit a knee-like shape with a break (or pivot) around the low- and high-mass halo limit. This feature arises due to AGN feedback, which redistributes baryons to the outskirts of low-mass haloes, reducing their baryon fractions and breaking self-similarity in these haloes (Ayromlou et al., 2023; Pop et al., 2022a). For this reason, the SPL model is insufficient for studying scaling relations across a wide range of haloes in the context of self-similarity (Pop et al., 2022a). To address this limitation, we require models capable of identifying the breakpoints in baryonic scaling relations, indicating where self-similarity is preserved. The models employed in this section are the BPL and GDPL models.
@@
The method we follow is similar to that used by P22. We perform a non-linear least squares regression to fit the SPL, BPL, and GDPL models to the median baryon fraction values in the IllustrisTNG (TNG100 and TNG300) samples. To compute the median values, we use 20 bins. This is the same bins used by P22, but in a higher mass range ($M_{500\mathrm{c}}\in[10^{12}-2\times 10^{15}]\,\mathrm{M}_{\odot}$). From these fits, we obtain the power-law parameters $\alpha_{\rm SPL}$, $\alpha_{\rm BPL,1}$, $\alpha_{\rm BPL,2}$, $\alpha_{\rm GDPL,1}$, and $\alpha_{\rm GDPL,2}$, along with the $\delta$ and $M_{\rm pivot}$ parameters from the GDPL model (see details in Equations 7, 8, and 9). The parameter $M_{\rm pivot}$ is a breakpoint, indicating whether self-similarity is maintained or broken.
@@
Figure 5: The Simple Power Law (SPL) fit (green region), the Broken Power Law (BPL) fit (magenta region) and the General Double Power Law (GDPL) fit (red region) of the IllustrisTNG haloes (median of the data in black solid line), to the scaling relation of the mass fraction of baryons $f_{\rm bar}$ and halo mass $M$ in a radius $R_{\rm 500c}$. All panels have their respective simulation tag in addition to a vertical dark green dashed line that represents the $M_{\rm pivot}$ of the broken fits. The power values of $\alpha_{\rm SPL}$ SPL fits, $\alpha_{\rm BPL,1}$ and $\alpha_{\rm BPL,2}$ BLP fits, and $\alpha_{\rm GDPL,1}$ and $\alpha_{\rm GDPL,2}$ GDPL fits, with the GDPL smoothness parameter $\delta$ and $M_{\rm pivot}$ value are in the Table 5. (Top panels) The fraction of baryons includes the hot-gas and the $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ aperture. (Bottom panels) The fraction of baryons includes all the gas and all stars inside the halo. The yellow region represent the Planck Collaboration et al. (2016a) fraction of baryons. All the regions are in $3\sigma$.
@@
Comparision type hot-gas and $r_{\rm op}$ stars all-gas and all stars
@@
Fig. 5 shows the SPL, BPL, and GDPL fits to the median values of the TNG100 and TNG300 halo samples, including the $M_{\rm pivot}$ value and the baryonic mass fraction of the Universe (Planck Collaboration et al., 2016a). All numerical values of the power-law parameters—$\alpha_{\rm SPL}$, $\alpha_{\rm BPL,1}$, $\alpha_{\rm BPL,2}$, $\alpha_{\rm GDPL,1}$, and $\alpha_{\rm GDPL,2}$—as well as the GDPL smoothness parameter $\delta$, the $M_{\rm pivot}$ value, and the respective model $\chi^{2}$-test results are presented in Table 5. The mean values and their associated errors are obtained from the fits to the median values of the sample. To compute the $\chi^{2}$-test, we use the median values of the TNG100 and TNG300 halo samples along with the models fitted with their mean parameter values. As shown in Fig. 5, the BPL and GDPL models exhibit a shape closely matching the distribution of IllustrisTNG haloes in the hot-gas + $r_{\rm op}$ case. In contrast, the SPL model matches the distribution for intermediate masses well but overestimates the baryon fraction at both low- and high-mass extremes. In the all-gas + all-stars case, all three models demonstrate similar shapes compared to the halo median values, although the SPL model slightly overestimates the baryon fraction at the low- and high-mass extremes, albeit less prominently than in the hot-gas + $r_{\rm op}$ case. The $\chi^{2}$-test results for the SPL model support these claims, showing significant differences compared to the BPL and GDPL models in the hot-gas + $r_{\rm op}$ case, but no notable differences in the all-gas + all-stars case. As detailed in Table 5, the $\chi^{2}$ values for the BPL and GDPL models are very close, with differences of $\chi^{2}_{\rm GDPL}-\chi^{2}_{\rm BPL}\in[0,0.001]$, which is effectively negligible for both models.
@@
In Table 5, we observe that the SPL parameter fits for TNG100 in both cases are closer to the parameters of the scaling relations in Fig. 3. However, the SPL parameter fits for TNG300 in both cases are less consistent with these scaling relation parameters. This discrepancy is not an issue, as a different fitting method is used in this analysis. Additionally, since both simulations use the same number of bins, TNG300 has a larger binning mass range than TNG100 due to the broader mass range of haloes in TNG300. We also find that the mean power-law parameters ($\alpha$) for the BPL and GDPL models are identical. The differences lie in their standard deviation values, where the BPL model shows smaller deviations compared to the GDPL model. As seen in Fig. 5, these standard deviations are negligible within the IllustrisTNG halo mass range but become notable outside this range. This effect is more apparent in TNG100 than in TNG300, as the median mass values in TNG300 exhibit a flatter shape, leading the BPL and GDPL models to fit similarly. The $M_{\rm pivot}$ values in the BPL and GDPL models are identical, with a range of mean values $\log_{10}\left(M_{\rm pivot}/{\rm M_{\odot}}\right)=13.80-14.07$. This range includes the $\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\sim 14$ mass limit reported by P22 as the threshold for the loss of self-similarity in scaling relations, and by A23 as the upper limit for baryon redistribution driven by AGN feedback. This strongly suggests that AGN feedback lowers the baryonic mass fraction in low-mass haloes. Despite considering all baryonic components within $R_{\rm 500c}$, low-mass haloes still have reduced baryon fractions, while high-mass haloes approach the baryonic fraction value reported by Planck Collaboration et al. (2016a) within $3\sigma_{\rm Planck}$. This supports the hypothesis that warm gas, cold gas, and ICL are underrepresented in observations. Finally, the mean softened transition parameter $\delta$ for the GDPL models is consistent across both simulations and comparison types, with $\delta\sim 0.25$. The $\sigma_{\delta}$ values are not significant, as $\delta$ primarily governs the transition between $\alpha_{1}$ and $\alpha_{2}$.
@@
The BPL and GDPL scaling relations could significantly aid observations in characterizing non-gravitational processes using the baryon fraction of haloes, particularly in the context of galaxy clusters. The fact that IllustrisTNG matches the SPL model slopes derived by C18 establishes a solid framework for analyzing similar observations and comparing, for instance, slope results with simulations. In the observational domain, ongoing and future surveys such as the Dark Energy Survey (Collaboration: et al., 2016), Euclid (Laureijs et al., 2011), the Vera C. Rubin Observatory’s LSST (LSST Dark Energy Science Collaboration, 2012), eROSITA (Merloni et al., 2012; Pillepich et al., 2012, 2018; Predehl et al., 2021; Bulbul et al., 2022), Athena X-ray observatory (Nandra et al., 2013), SPT-3G (Benson et al., 2014; Bender et al., 2018), CMB-S4 (Abazajian et al., 2016), CMB-HD (Sehgal et al., 2019), and Advanced ACTpol (Henderson et al., 2016) could validate and benefit from these models by comparing them with current and future cosmological hydrodynamical simulations. For instance, upcoming projects such as TNG-Cluster simulations (Nelson et al., 2024; Ayromlou et al., 2024; Lee et al., 2024; Lehle et al., 2024; Rohr et al., 2024; Truong et al., 2024), the MillenniumTNG Project (Hernández-Aguayo et al., 2023; Pakmor et al., 2023; Barrera et al., 2023; Kannan et al., 2023; Hadzhiyska et al., 2023, 2023a, 2023b; Bose et al., 2023; Contreras et al., 2023; Delgado et al., 2023; Ferlito et al., 2023), and the FLAMINGO simulations (Schaye et al., 2023; Kugel et al., 2023) could provide valuable insights. These simulations will also allow the extension of this framework by incorporating a larger sample of haloes, enabling a deeper understanding of baryonic scaling relations and their implications for cosmology.

### v2

@@ 4.2 Looking for the missing baryons
A common limitation in observations is the lack of information on the cold and warm gas components. As simulations do not have this limitation, we estimated the fraction of baryons that remain undetected when only the hot gas in the ICM is considered. We divided the baryonic matter into cold, warm, and hot gas for the gaseous component, and into galaxy stellar mass (from $r_{\rm op}$ and with Subfind) and the ICL for the stellar component. We further separated our clusters into low-mass $\log_{10}{(M_{\rm 500c}/{\rm M_{\odot}})}<14$ and high-mass haloes $\log_{10}{(M_{\rm 500c}/{\rm M_{\odot}})}\geq 14$. Our sample comprises 137 and 1115 low-mass haloes in IllustrisTNG-100 and IllustrisTNG-300, and 81 and 490 high-mass haloes in IllustrisTNG-100 and IllustrisTNG-300, respectively. For visualisation purposes, we also include the complete sample (labelled ”all halo types”) in the following figures.
@@
Furthermore, Table 3 shows the average baryonic mass values for components not considered in the C18 observations: cold gas, warm gas, and ICL ($r_{\rm op}$). We defined the sum of those components as the missing baryons in simulations. The data in Table 3 reveal a tendency for low-mass haloes to contain more missing (undetected) baryons than high-mass haloes. Specifically, low-mass haloes exhibit missing baryon fractions of approximately $17.83$ per cent in TNG100 and $17.67$ per cent in TNG300. In contrast, high-mass haloes have missing baryon fractions of about $6.95$ per cent in TNG100 and $5.90$ per cent in TNG300. In both cases, hot gas remains the most significant baryonic component within the haloes. However, warm gas and ICL ($r_{\rm op}$) correspond to a non-negligible fraction of baryons missed in low-mass haloes. In high-mass haloes, most of the missing baryons are attributed to the ICL ($r_{\rm op}$), which shows a similar proportional contribution to that observed in low-mass haloes. Finally, Table 3 also presents the stellar components measured using the Subfind aperture, highlighting the differences between this method and the $r_{\rm op}$ aperture in terms of stellar mass measurements. The Subfind algorithm collects more stellar mass within galaxies compared to the $r_{\rm op}$ spherical aperture, as it assigns all particles within the virial radius that are not associated with satellite subhaloes to the central galaxy (Springel et al., 2001; Dolag et al., 2009). As a result, diffuse stellar components, such as the stellar halo, are included as part of the central galaxy mass.
When considering all haloes, Table 3 shows that TNG100 and TNG300 have missing baryon fractions of 13.79 per cent and 14.08 per cent, respectively. Hot gas remains the most significant baryonic component. The overall fractions are similar to those of low-mass haloes but with lower contributions from warm gas and the ICL ($r_{\rm op}$) components, as high-mass haloes are included in this analysis. Warm gas is the most affected component when considering all haloes, showing a reduction of approximately three per cent compared to low-mass haloes in both simulations. The ICL ($r_{\rm op}$) component also exhibits a reduction of about $0.40$ per cent relative to the low-mass haloes in both simulations. This analysis highlights the importance of dividing the halo sample, as the baryonic content can differ between high- and low-mass clusters.
@@
Gouin et al. (2022) studied the distribution of gas components in TNG300 haloes at $z=0$. They categorised these haloes by mass into clusters ($M_{\rm 200c}/{\rm{M_{\odot}}}/h>1\times 10^{14}$) and groups ($5\times 10^{13}<M_{\rm 200c}/{\rm{M_{\odot}}}/h<1\times 10^{14}$) of galaxies. Additionally, the gas was divided into five independent phases: hot gas, warm dense gas (WCGM), warm diffuse gas (WHIM), cold dense gas (halo gas), and cold diffuse gas (IGM). The gas fraction of each phase was calculated as a function of $\rm R_{200}$. Table 4 presents the relative gas percentages for our haloes, calculated from the percentages in Table 3. Despite the differences between our sample and that of Gouin et al. (2022), the gas fractions are remarkably similar at $R\sim 0.6R_{200}\sim R_{500}$. In galaxy groups, most of the contribution to these percentages, apart from hot gas, comes from warm diffuse gas, which corresponds to the circumgalactic medium (CGM) (Gouin et al., 2022). At this radius, WCGM gas accounts for a larger fraction than WHIM gas. In galaxy clusters, the majority of the gas fraction comes from hot gas. This analysis suggests that the gas phase associated with the missing baryons in our sample is likely dominated by WCGM gas.
These results suggest that the break in the power law seen in Fig. 3 is driven by the higher fractions of ICL, cold gas, and warm gas present in low-mass clusters compared to massive ones. The percentages of cold gas are lower in all cases compared to those of warm gas. Therefore, warm gas and the ICL in low-mass haloes are the primary contributors to the observed changes in baryonic scaling relations. This interpretation aligns with the WHIM proposal, which suggests that the missing baryons identified in C18 could reside in the warm diffuse phase. Furthermore, A23 propose that missing baryons are redistributed by non-gravitational processes (e.g. AGN, stellar, and UV feedback) into a region between $1.5-2.5R_{\rm 200c}$, where the baryon fraction reaches the universal value. Both scenarios are compatible and suggest a coherent explanation for the missing baryons problem. These findings may have significant implications for future observational research related to the baryon distribution, mass measurements, and the calibration of scaling relations.
@@ 4.3 FoF masses in baryonic scaling relations
We explored changes in baryonic scaling relations when considering the total mass of haloes within a sphere radius $R_{\rm 500c}$, as discussed in Subsections 4.1 and 4.2. The $R_{\rm 500c}$ was particularly useful for measuring gas masses in observations, especially those using X-ray and the SZ effect. In this section, we switch from using masses defined within $R_{\rm 500c}$ to those defined by the FoF algorithm. This approach accounts for all the mass bound to the cluster without a fixed aperture and allowed us to compare how scaling relations change when using these broader mass definitions.
@@
Figure 4: SPL fit (green region) of the IllustrisTNG haloes (black dots), to the scaling relation of the mass fraction of baryons $f_{\rm bar}$ and halo mass $M$ measured by the FoF algorithm, for all redshifts. Each panel is labelled with the corresponding simulation tag and the value of the $\alpha_{\rm SPL}$ slope in the upper-left corner. The baryon fraction includes all gas and stellar mass within each halo. The yellow region represents the baryon fraction from Planck Collaboration et al. (2016a) All regions are in $3\sigma$.
@@
Fig. 4 presents the SPL scaling relation models for baryonic mass fractions of IllustrisTNG haloes in the FoF-delimited mass context. Comparing these results with the all-gas + all-stars scaling relations at $R_{\rm 500c}$ shown in Fig. 3, we observed a mean slope difference of $\alpha_{\rm SPL,FoF}-\alpha_{\rm SPL,500c}\simeq-0.06$ in TNG100 and $\alpha_{\rm SPL,FoF}-\alpha_{\rm SPL,500c}\simeq 0.08$ in TNG300. In the FoF context, low-mass haloes increase their baryon fractions, displaying a tendency towards $\alpha_{\rm SPL,FoF}\to 0$ slopes in both simulations. This behaviour accounts for the differences observed with the $R_{\rm 500c}$ mass context. Furthermore, the self-similarity of the baryon fraction is satisfied, as the $\alpha_{\rm SPL,FoF}\to 0$ results are consistent with Eq. (6). Furthermore, these slope tendencies indicate that high-mass haloes can form through the assembly of low-mass haloes, reflecting the hierarchical formation scenario. However, most IllustrisTNG haloes do not reach the baryon fraction measured by Planck Collaboration et al. (2016a) within the $3\sigma_{\rm Planck}$ region (with $\sigma_{\rm Planck}=0.004$ being the error of the baryon fraction in Planck Collaboration et al., 2016a). Most haloes exhibit baryon fractions approximately $0.1$ dex below the mean value, with only a few exceptions. This result suggests the re-emergence of the missing baryon problem in the FoF context, which will be further discussed in Section 5.
4.4 Fitting power-law models as scaling relations
@@
Figure 3, shows the SPL model fits to the scaling relations of baryon fractions in IllustrisTNG haloes. This provides a framework for comparison with the scaling relations from C18. However, the distribution of scattered data and the SPL model fit show significant deviations. In the hot gas + $r_{\rm op}$ star comparison, the data exhibit a knee-like shape with a break (or pivot) around the low- and high-mass halo limit. This feature arises due to AGN feedback, which redistributes baryons to the outskirts of low-mass haloes, reducing their baryon fractions and breaking self-similarity in these haloes (Ayromlou et al., 2023; Pop et al., 2022a). As a result, the SPL model is insufficient for studying scaling relations across a wide range of haloes in the context of self-similarity (Pop et al., 2022a). To address this limitation, models capable of identifying breakpoints in the baryonic scaling relations are needed to determine where self-similarity is preserved. In this section, we employ the BPL and GDPL models for this purpose.
@@
The method we followed was similar to that used by P22. We performed a non-linear least squares regression to fit the SPL, BPL, and GDPL models to the median baryon fraction values in the IllustrisTNG (TNG100 and TNG300) samples. To compute the median values, we used 20 bins, as in P22, but applied them to a higher mass range ($M_{500\mathrm{c}}\in[10^{12}-2\times 10^{15}]\,\mathrm{M}_{\odot}$). From these fits, we obtained the power-law parameters $\alpha_{\rm SPL}$, $\alpha_{\rm BPL,1}$, $\alpha_{\rm BPL,2}$, $\alpha_{\rm GDPL,1}$, and $\alpha_{\rm GDPL,2}$, along with the $\delta$ and $M_{\rm pivot}$ parameters from the GDPL model (see details in Equations 7, 8, and 9). The parameter $M_{\rm pivot}$ represents a breakpoint, indicating whether self-similarity was maintained or broken.
@@
Figure 5: SPL fit (green region), BPL fit (magenta region) and GDPL fit (red region) of the IllustrisTNG haloes (median of the data in black solid line), to the scaling relation of the mass fraction of baryons $f_{\rm bar}$ and halo mass $M$ in a radius $R_{\rm 500c}$. Each panel is labelled with the corresponding simulation tag in addition to a vertical dark green dashed line that represents the $M_{\rm pivot}$ of the broken fits. The power values of $\alpha_{\rm SPL}$ SPL fits, $\alpha_{\rm BPL,1}$ and $\alpha_{\rm BPL,2}$ BLP fits, and $\alpha_{\rm GDPL,1}$ and $\alpha_{\rm GDPL,2}$ GDPL fits, with the GDPL smoothness parameter $\delta$ and $M_{\rm pivot}$ value are in the Table 5. (Top panels) The fraction of baryons includes the hot-gas and the $2\times R_{M_{\rm star}/2}\sim r_{\rm op}$ aperture. (Bottom panels) The baryon fraction includes all gas and stellar mass within each halo. The yellow region represent the Planck Collaboration et al. (2016a) fraction of baryons. All the regions are in $3\sigma$.
@@
Comparison type hot-gas and $r_{\rm op}$ stars all-gas and all stars
@@
Figure 5 shows the SPL, BPL, and GDPL fits to the median values of the TNG100 and TNG300 halo samples, including the $M_{\rm pivot}$ value and the baryonic mass fraction of the Universe (Planck Collaboration et al., 2016a). All numerical values of the power-law parameters—$\alpha_{\rm SPL}$, $\alpha_{\rm BPL,1}$, $\alpha_{\rm BPL,2}$, $\alpha_{\rm GDPL,1}$, and $\alpha_{\rm GDPL,2}$—as well as the GDPL smoothness parameter $\delta$, the $M_{\rm pivot}$ value, and the respective model $\chi^{2}$-test results are presented in Table 5. The mean values and their associated errors were obtained from the fits to the median values of the sample. To compute the $\chi^{2}$-test, we used the median values of the TNG100 and TNG300 halo samples along with the models fitted with their mean parameter values. As shown in Fig. 5, the BPL and GDPL models exhibit a shape that closely matches the distribution of IllustrisTNG haloes in the hot-gas + $r_{\rm op}$ case. In contrast, the SPL model matches the distribution for intermediate masses well but overestimates the baryon fraction at both low- and high-mass extremes. In the case of all gas + all stars, all three models demonstrate similar shapes compared to the halo median values, although the SPL model slightly overestimates the baryon fraction at the low- and high-mass extremes, but this effect is less pronounced than in the hot-gas + $r_{\rm op}$ case. The $\chi^{2}$-test results for the SPL model support these claims, showing significant differences compared to the BPL and GDPL models in the hot-gas + $r_{\rm op}$ case, but no notable differences in the all-gas + all-stars case. As detailed in Table 5, the $\chi^{2}$ values for the BPL and GDPL models were very close, with differences of $\chi^{2}_{\rm GDPL}-\chi^{2}_{\rm BPL}\in[0,0.001]$, which is effectively negligible for both models.
@@
In Table 5, we find that the SPL parameter fits for TNG100 in both cases are closer to the parameters of the scaling relations in Fig. 3. However, the SPL parameter fits for TNG300 in both cases are less consistent with these scaling relation parameters. This discrepancy is due to the different fitting method used in this analysis. Since both simulations use the same number of bins, TNG300 has a larger binning mass range than TNG100 due to its broader halo mass range. We also find that the mean power-law parameters ($\alpha$) for the BPL and GDPL models are identical, but the BPL model shows smaller deviations than the GDPL model. As shown in Fig. 5, these standard deviations are negligible within the IllustrisTNG halo mass range but become noticeable outside of this range. This effect is more apparent in TNG100 than in TNG300, as the median mass values in TNG300 exhibit a flatter shape, resulting in similar fits for the BPL and GDPL models. The $M_{\rm pivot}$ values in the BPL and GDPL models are identical, with mean values in the range $\log_{10}\left(M_{\rm pivot}/{\rm M_{\odot}}\right)=13.80-14.07$. This range includes the $\log_{10}\left(M_{\rm 500c}/{\rm M_{\odot}}\right)\sim 14$ mass limit reported by P22 as the threshold for the loss of self-similarity in scaling relations, and by A23 as the upper limit for baryon redistribution driven by AGN feedback. This strongly suggests that AGN feedback lowers the baryonic mass fraction in low-mass haloes. Despite considering all baryonic components within $R_{\rm 500c}$, low-mass haloes still show reduced baryon fractions, while high-mass haloes approach the baryonic fraction value reported by Planck Collaboration et al. (2016a) within $3\sigma_{\rm Planck}$. This supports the hypothesis that warm gas, cold gas, and ICL are under-represented in observations. Finally, the mean softened transition parameter, $\delta$, for the GDPL models was consistent across both simulations and comparison types, with $\delta\sim 0.25$. The $\sigma_{\delta}$ values are not significant, as $\delta$ primarily governs the transition between $\alpha_{1}$ and $\alpha_{2}$.
@@
The BPL and GDPL scaling relations provide improved characterisation of non-gravitational processes using the baryon fraction of haloes, particularly in the context of galaxy clusters. The SPL model slopes derived from IllustrisTNG are consistent with those reported by C18, supporting the analysis of similar observations and the comparison of slope results with simulations. In the observational domain, ongoing and future surveys such as the Dark Energy Survey (Collaboration: et al., 2016), Euclid (Laureijs et al., 2011), the Vera C. Rubin Observatory’s LSST (LSST Dark Energy Science Collaboration, 2012), eROSITA (Merloni et al., 2012; Pillepich et al., 2012, 2018; Predehl et al., 2021; Bulbul et al., 2022), Athena X-ray observatory (Nandra et al., 2013), SPT-3G (Benson et al., 2014; Bender et al., 2018), CMB-S4 (Abazajian et al., 2016), CMB-HD (Sehgal et al., 2019), and Advanced ACTpol (Henderson et al., 2016) could validate and benefit from these models by comparing them with current and future cosmological hydrodynamical simulations. For instance, upcoming projects such as TNG-Cluster simulations (Nelson et al., 2024; Ayromlou et al., 2024; Lee et al., 2024; Lehle et al., 2024; Rohr et al., 2024; Truong et al., 2024), the MillenniumTNG Project (Hernández-Aguayo et al., 2023; Pakmor et al., 2023; Barrera et al., 2023; Kannan et al., 2023; Hadzhiyska et al., 2023, 2023a, 2023b; Bose et al., 2023; Contreras et al., 2023; Delgado et al., 2023; Ferlito et al., 2023), and the FLAMINGO simulations (Schaye et al., 2023; Kugel et al., 2023) could provide valuable insights. These simulations will also allow for the extension of this framework by incorporating a larger sample of haloes, enabling a deeper understanding of baryonic scaling relations and their implications for cosmology.